OpenStudy (anonymous):

A conference room is in the shape of a rectangle. Its floor has a length of (x − 4) meters and a width of (3x − 1) meters. The expression below represents the area of the floor of the room in square meters. (x − 4)(3x − 1) Which of the following simplified expressions represents the area of the floor of the conference room in square meters? Can you walk me through it?

OpenStudy (unklerhaukus):

\[(x − 4)(3x − 1)=x (3x − 1)− 4(3x − 1)\\\qquad\qquad\qquad=x(3x)+x(-1)-4\times(3x)-4(-1)\\\qquad\qquad\qquad=\]

OpenStudy (anonymous):

3x^2-x-4*3x+4= -x*7=-7x?

OpenStudy (unklerhaukus):

you've got this bit right \[=3x^2-x-(4\times3)x+4\\=\]

OpenStudy (anonymous):

oh I didnt know the 4 and 3 were in perenthisies

OpenStudy (anonymous):

3x-12x^2+4?

OpenStudy (unklerhaukus):

\[=3x^2-x-12x+4\\=3x^2-(1+12)x+4\\=\]

OpenStudy (anonymous):

3x^2-13x+4?

OpenStudy (unklerhaukus):

yeah , thats it !

OpenStudy (anonymous):

Oh thank you! Could you help me with one more?

OpenStudy (unklerhaukus):

sure

OpenStudy (anonymous):

A glass cylinder is completely filled with 32m6n 2 cubic inches of salt solution. The jug has a base area of 8m3n 2 square inches. The height of the jug, in inches, is represented by the expression 32m^6n^2/8m^3n^2. Which of the following simplified expressions represents the height of the jug in inches?

OpenStudy (unklerhaukus):

\[\frac{32m^6n^2}{8m^3n^2}\]

OpenStudy (unklerhaukus):

\[\frac{x^ny^m}{x^py^q}=x^{n-p}y^{m-q}\]

OpenStudy (anonymous):

I cant read whats by the x and y

OpenStudy (unklerhaukus):

\[\Large\frac{x^ny^m}{x^py^q}=x^{n-p}y^{m-q}\] can you read it now?

OpenStudy (anonymous):

yes sorry about that.

OpenStudy (anonymous):

but what do I do with that?

OpenStudy (unklerhaukus):

well in your case \[\frac{32m^6n^2}{8m^3n^2}=\frac{32}8m^{6-3}n^{2-2}=\]

OpenStudy (anonymous):

@UnkleRhaukus Sorry my computer went all bonkers. Can you still help me?

OpenStudy (unklerhaukus):

can you see the first step i have made? (6-3)= now you can simplify those indicies...